Find the confidence interval given the following data Cl-95, s-3, n-15, ×-25





1
Expert's answer
2022-03-29T08:34:09-0400

"\left( \bar{x}-\frac{s}{\sqrt{n}}t_{n-1,\frac{1+\gamma}{2}},\bar{x}+\frac{s}{\sqrt{n}}t_{n-1,\frac{1+\gamma}{2}} \right) =\\=\left( 25-\frac{3}{\sqrt{15}}\cdot 2.1448,25+\frac{3}{\sqrt{15}}\cdot 2.1448 \right) =\\=\left( 23.3386,26.6614 \right)"